OpenStudy (anonymous):

A square field is to be hydro-seeded. The field has a diagonal length of 225 yards. How many square feet need to be hydro-seeded?

OpenStudy (cwrw238):

first find area in square yds: 225^2 = 2x^2 x^2 = 25312.5 = 25312.5 / 9 = 2812.5 sq yds

OpenStudy (anonymous):

thnx

OpenStudy (cwrw238):

yw